cbagwe8982:

Using only addition, how do you add eight 8's and get the number 1000?

jozelynns:

1000 = x + 992 where x = 8

Ashtiel:

1000 divided by 8 = 125

AZ:

888 + 88 + 8 + 8 + 8 = 1000